A charge of 4.15 mC is placed at each corner of a square 0.310 m on a side.

1. Determine the magnitude of the force on each charge.
Answer #1

Since all corners have same charge therefore if we find force at one charge then all others charge will also experience the same force.
Now force on charge 4 are shown in the figure due to other charges.
Force on 4 due to 1
F_{41} = rac{kq_{1}*q_{4}}{a^{2}}
On putting the values we get
9 10 4.151015 10 41 - 0.312
F41 = 1.613*106 N
Now force on charge 4 due to charge 3
F_{43} = rac{kq_{3}*q_{4}}{a^{2}}
9 10 4.151015 10 43 0.312
F43 = 1.613*106 N
Now force due to charge 2
kg2 (av2)2 42 =
9 10 4.151015 10 F42 = (0.31V2)2
F42 = 8.06*105 N
Now the force in X direction
FX = F41 + F42Cos45 = 1.613*106 + (8.06*105)Cos45 = 21.83*105 N
In the Y direction
FY = F43 + F42Sin45 = 1.613*106 + (8.06*105)Cos45 = 21.83*105 N
Hence the magnitude of the force will be
3 2
F = 30.87*105 N
Similarly force on all charges will be 30.87*105 N due to symmetry.
F42 43 a-0.31 m F4.15 mC* 9415 mC 41 0.31 45 7415 mC q,-4.15 mC